OpenStudy (anonymous):

Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. Point A: (1,4) Point B: (3,5)

hero (hero):

Point Slope form: \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

Yeah, I know that part, I'm just confused about where to plug the numbers in

hero (hero):

Only \(x_1, y_1\) and \(m\) need to be filled in

hero (hero):

\((x_1,y_1) = (1,4)\)

hero (hero):

\(m = \dfrac{y_2 - y_1}{x_2 - x_1}\)

OpenStudy (anonymous):

oh okay, and m is the slope, but i'm not sure how to fund that

OpenStudy (anonymous):

*find

hero (hero):

\((x_2, y_2) = (3,5)\)

hero (hero):

\(m = \dfrac{5 - 4}{3 - 1}\)

OpenStudy (anonymous):

so 1 over 2

hero (hero):

\(m = \dfrac{1}{2}\) So now you have everything you need to plug-in

OpenStudy (anonymous):

okay let me see if this is right 5y-4y=1/2(3-1)?

OpenStudy (anonymous):

Hope I'm right! Thanks for the help! ^-^